# Homogeneous $\mathrm{G}_{2}$ and Sasakian instantons on the Stiefel $7$-manifold

### v1

@@ 4.3 Stability of SDCI and $\mathrm{G}_{2}$-instantons
Let $E$ be a Sasakian ${\rm G}$-bundle over a compact closed 7-manifold. For any variation $\nabla^{\prime}=\nabla+\alpha\in\mathcal{A}(E)$, $\alpha\in\Omega^{1}({\mathfrak{g}}_{E})$, on the underlying complex vector bundle E, we have
$\kappa(\nabla)=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\ast(\omega\otimes s)),$ (4.22)
where $s\in\Gamma(E)_{B}$ is some transversal section of $E$.
@@
The proof of [portilla2023, Lemma 3.14] provide that for an arbitrary deformation $\nabla^{\prime}=\nabla+\alpha$, where $\alpha\in\Omega^{1}({\mathfrak{g}}_{E})=\Omega^{1}_{H}({\mathfrak{g}}_{E})\oplus\eta\wedge\Omega^{0}_{H}({\mathfrak{g}}_{E})$, the charge varies as
$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\alpha)\wedge\omega^{2}$ (4.23)
$\quad+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(d_{\nabla}\alpha\wedge\alpha)\wedge\omega^{2}+\frac{1}{3}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\alpha\wedge\alpha\wedge\alpha)\wedge\omega^{2}.$
In particular, assuming a transversal deformation, (4.21) holds because the defect term in (4.23) vanishes, as the subintegral quantities in the last two terms belong to $\Omega^{7}_{H}({\mathfrak{g}}_{E})=0$. Hence, we can assume that $\alpha$ is a vertical $1$-form, i.e., $\alpha=\eta\otimes s\in\eta\wedge\Omega^{0}_{H}({\mathfrak{g}}_{E})$ for some transversal section $s$, therefore
$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge(\eta\otimes s))\wedge\omega^{2}+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(d_{\nabla}(\eta\otimes s)\wedge(\eta\otimes s))\wedge\omega^{2}$
$=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\eta\wedge\omega^{2}\otimes s)+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\omega\otimes s-\eta\otimes\nabla s)\wedge(\eta\otimes s))\wedge\omega^{2}$
$=\kappa(\nabla^{\prime})+\int_{M}\mathop{\mathrm{Tr}}\nolimits({\rm F}_{\nabla}\wedge\ast(\omega\otimes s))+\frac{1}{2}\int_{M}\mathop{\mathrm{Tr}}\nolimits(\eta\wedge\omega^{3}\otimes[s,s]),$
hence, (4.22) follows. ∎
@@
Let $E$ be a Sasakian ${\rm G}$-bundle over a compact closed 7-manifold. Any SDCI is a local minimizer of the YM-functional (4.14).
@@
From (4.19), note that the YM-functional assumes the value of the charge $\kappa(\nabla)$ (4.17). From (4.22)
$\kappa(\nabla)$ $=\kappa(\nabla^{\prime})+({\rm F}_{\nabla},\omega\otimes s)_{L^{2}}$
$=\kappa(\nabla^{\prime}),$
the last line follows since $\Omega^{2}_{8}({\mathfrak{g}}_{E})\perp\omega\otimes\Omega^{0}({\mathfrak{g}}_{E})$. Hence, if $\nabla$ is a SDCI the charge is locally constant. The same is true if $\nabla$ is a ASDCI, since $\Omega^{2}_{6}({\mathfrak{g}}_{E})\perp\omega\otimes\Omega^{0}({\mathfrak{g}}_{E})$. ∎

### v2

@@ 5.1 Infinitesimal deformation of $\mathrm{G}_{2}$-instantons
Let $\zeta$ be the spinor associated with $\varphi$, then it holds
$(X\lrcorner\varphi)\cdot\zeta=3X\cdot\zeta,\qquad(X\lrcorner\gamma)\cdot\zeta=2h(X)\cdot\zeta\quad\text{and}\quad(X\lrcorner\ast\gamma)\cdot\zeta=-2h(X)\cdot\zeta,$
where $X\in\mathscr{X}(M)$, $\gamma\in\Omega^{3}_{27}$ and $h$ is the traceless symmetric tensor defining $\gamma$ by
$\gamma_{ijk}=h_{ip}\gamma_{pjk}+h_{jp}\gamma_{ipk}+h_{kp}\gamma_{ijp}.$
@@
Now, using (5.2), we get
[…]
$=3g_{im}\,(0,e^{m})=3\,(0,e_{i})=3e_{i}(0,1)$
$=3e_{i}\cdot\zeta$
Here, we used the identity $\varphi_{ijk}\varphi_{mnk}=g_{im}g_{jn}-g_{in}g_{jm}+\psi_{ijmn}$ [karigiannis2009flows]*Lemma A.8. Assume $\gamma\in\Lambda^{3}_{27}$, we compute
$(X\;\lrcorner\;\gamma)\cdot\zeta$ $=X_{i}\,(e_{i}\;\lrcorner\;\gamma)\cdot\zeta=\frac{1}{2}\,X_{i}\,\gamma_{ijk}\,e^{j}\cdot e^{k}\cdot\zeta$
$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+h_{j}{}^{p}\,\varphi_{ipk}+h_{k}{}^{p}\,\varphi_{ijp}\bigr)e^{j}\cdot e^{k}\cdot\zeta.$
$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+2h_{j}{}^{p}\,\varphi_{ipk}\bigr)e^{j}\cdot e^{k}\cdot(1,0)$
$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+2h_{j}{}^{p}\,\varphi_{ipk}\bigr)(0,\,e^{j}\times e^{k})$
$=\frac{1}{2}\,X_{i}\bigl(h_{i}{}^{p}\,\varphi_{pjk}+2h_{j}{}^{p}\,\varphi_{ipk}\bigr)(0,\,\varphi^{jk}{}_{m}e^{m})$
$=\frac{1}{2}\,X_{i}\,h_{i}{}^{p}\,(6\,g_{pm})\,(0,e^{m})+X_{i}\,h_{j}{}^{p}\bigl(g_{im}g^{j}{}_{p}-g_{i}{}^{j}g_{pm}+\psi_{ipm}{}^{j}\bigr)(0,e^{m})$
$=3\,h(X)\cdot\zeta-h(X)\cdot\zeta.$
$=2\,h(X)\cdot\zeta$
The next result characterises the infinitesimal deformations of a $\mathrm{G}_{2}$-instanton in terms of the operator (5.7).
Proposition 5.5.
Let $\varepsilon\in\Omega^{1}(M,\mathrm{Ad}(P))$ and $D^{t,A}$ the Dirac operator (5.7) with $A$ a $\mathrm{G}_{2}$-instanton. Then $\varepsilon$ defines an infinitesimal deformation of AA (cf. Definition 5.2), if and only if
$D^{t,A}(\varepsilon\cdot\zeta)=\frac{(-5+t)}{8}\,\tau_{0}\,\varepsilon\cdot\zeta+(3t-1)\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
Proof.
Let $\{e_{i}\}_{i=1}^{7}$ be a local orthonormal frame on $M$. By definition of the twisted Dirac operator associated with the product connection $\nabla^{0,A}$, a direct computation gives
$D^{0,A}(\varepsilon\cdot\zeta)$ $=\sum e_{i}\cdot\nabla^{0,A}_{e_{i}}(\varepsilon\cdot\zeta)=(d_{A}\varepsilon+d_{A}^{\ast}\varepsilon)\cdot\zeta+\sum e_{i}\cdot\varepsilon\cdot\nabla^{0,A}_{e_{i}}(\zeta)$
$=(d_{A}\varepsilon+d_{A}^{*}\varepsilon)\cdot\zeta-\sum_{i}\left(\frac{\tau_{0}}{24}e_{i}\cdot\varepsilon\cdot(e_{i}\lrcorner\varphi)\cdot\zeta-\frac{1}{4}e_{i}\cdot\varepsilon\cdot(e_{i}\lrcorner\tau_{3})\cdot\zeta\right).$
Here, we used that $X\cdot Y+Y\cdot X=-2g(X,Y)1$ and $\nabla^{1}\zeta=0$ for the connection (5.5). Now, using Lemma 5.4 we get
$D^{0,A}(\varepsilon\cdot\zeta)$ $=(d_{A}\varepsilon+d_{A}^{*}\varepsilon)\cdot\zeta-\frac{5\tau_{0}}{8}\,\varepsilon\cdot\zeta-(\varepsilon\lrcorner\tau_{27})\cdot\zeta.$
On the other hand, for $H_{\varphi}=\frac{\tau_{0}}{6}\,\varphi-\tau_{3}$ we have
[…]
$=\frac{1}{2}\sum_{i,j,k}H_{ijk}\,\varepsilon_{\ell}\cdot e^{i}\cdot e^{\ell}\cdot e^{j}\cdot e^{k}\cdot\zeta+\sum H_{i\ell k}\varepsilon_{\ell}e^{i}\cdot e^{k}\cdot\zeta-2\,(\varepsilon\;\lrcorner\;H)\cdot\zeta$
$=-\frac{1}{2}\sum H_{ijk}\varepsilon_{\ell}\cdot e^{\ell}\cdot e^{i}\cdot e^{j}\cdot e^{k}\cdot\zeta-\sum H_{\ell jk}\,\varepsilon_{\ell}e^{j}\cdot e^{k}\cdot\zeta-4\,(\varepsilon\;\lrcorner\;H)\cdot\zeta.$
$=-3\,\varepsilon\cdot H\cdot\zeta-6\,(\varepsilon\;\lrcorner\;H)\cdot\zeta=-\frac{1}{2}\,\tau_{0}\,\varepsilon\cdot\varphi\cdot\zeta-6\,(\varepsilon\;\lrcorner\;H)\cdot\zeta.$
$=\frac{7}{2}\,\tau_{0}\,\varepsilon\cdot\zeta-\tau_{0}\,(\varepsilon\;\lrcorner\;\varphi)\cdot\zeta+6\,(\varepsilon\;\lrcorner\;\tau_{3})\cdot\zeta.$
$=\frac{1}{2}\,\tau_{0}\,\varepsilon\cdot\zeta+12\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
Consequently, from (5.7) we obtain
$D^{t,A}(\varepsilon\cdot\zeta)=(d_{A}\varepsilon+d_{A}^{*}\varepsilon)\cdot\zeta-\frac{5}{8}\,\tau_{0}\,\varepsilon\cdot\zeta-(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta+\frac{t}{8}\,\tau_{0}\,\varepsilon\cdot\zeta+3t\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
Assume that $\varepsilon$ satisfies the linearised instanton equation together with the Coulomb gauge condition, i.e., $d_{A}\varepsilon=0$ and $d_{A}^{*}\varepsilon=0$, then
$D^{t,A}(\varepsilon\cdot\zeta)=\frac{-5+t}{8}\,\tau_{0}\,\varepsilon\cdot\zeta+(3t-1)\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$ (5.8)
Since $\zeta$ is parallel with respect to $\nabla^{1}$, then $D^{1,A}(\varepsilon\cdot\zeta)=(D^{1,A}\varepsilon)\cdot\zeta,$ in consequence,
$(D^{1,A}\varepsilon)\cdot\zeta=-\frac{\tau_{0}}{2}\,\varepsilon\cdot\zeta+2\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$ (5.9)
Hence, the infinitesimal deformation space of a $\mathrm{G}_{2}$-instanton $A$ on a $7$-manifold with coclosed $\mathrm{G}_{2}$-structure is isomorphic to the kernel of the operator
$D^{1,A}+\frac{1}{2}\,\tau_{0}\,\mathrm{Id}_{\Omega^{1}(M,\mathrm{Ad}(P))}-2\,\tau_{27}^{*}\otimes\mathrm{Id}_{\mathrm{Ad}(P)}\in\mathrm{End}\bigl(\Omega^{1}(M,\mathrm{Ad}(P))\bigr).$ (5.10)
We have $D^{t,A}=D^{1,A}+(t-1)\,\frac{1}{4}\,g^{-1}\not${H}. Consequently,
$D^{t,A}(\varepsilon\cdot\zeta)=D^{1,A}(\varepsilon\cdot\zeta)+(t-1)\,\frac{\tau_{0}}{8}\,\varepsilon\cdot\zeta+(t-1)\,3\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta.$
Setting $t=\frac{1}{3}$, we obtain $D^{1/3,A}(\varepsilon\cdot\zeta)=D^{1,A}(\varepsilon\cdot\zeta)-\frac{\tau_{0}}{12}\,\varepsilon\cdot\zeta-2\,(\varepsilon\;\lrcorner\;\tau_{27})\cdot\zeta$ and consequently,
$D^{1/3,A}+\frac{7}{12}\,\tau_{0}\,\mathrm{Id}=D^{1,A}+\frac{1}{2}\,\tau_{0}\,\mathrm{Id}-2\,\tau_{27}^{*}\otimes\mathrm{Id}_{\mathrm{Ad}(P)}$ (5.11)
5.2 Weitzenböck inequality and curvature obstruction
It is known the Weitzenböck–Lichnerowicz formula with torsion cf. [Agricola2004Holonomy]
Proposition 5.6.
[…]
$(D^{t/3,A})^{2}\mu=(\nabla^{t,A})^{\ast}\nabla^{t,A}\mu+\frac{1}{4}\,\mathrm{Scal}\,\mu+\frac{t}{6}\,d\varphi\cdot\mu-\frac{t^{2}}{18}\,\|\varphi\|^{2}\mu+F\cdot\mu.$ (5.12)
As an application of Propositions 5.5 and 5.6, we prove the following rigidity result.
@@ 5.2 Weitzenböck inequality and curvature obstruction
Let $(V^{5,2},\varphi)$ be the Stiefel manifold endowed with the coclosed $\mathrm{G}_{2}$–structure given in Corollary 2.8 and let AA denote the $\mathrm{G}_{2}$–instanton (4.9) on the ${\rm SO}(3)$-bundle $P_{1}=V^{5,2}\times_{\phi_{1}}{\rm SO}(3)$. Then, for $y_{2}>0$ as in (4.12) there exists a non–empty interval $y_{2}\in I$ such that, for all values $y_{2}\in I$, the $\mathrm{G}_{2}$–instanton $A$ is rigid for these values of the metric parameter $y_{2}$.
@@
[…]
$\bigl|\nabla^{1/3,A}(\varepsilon\cdot\zeta)\bigr|^{2}$ $=\left\langle(D^{\frac{1}{3},A})^{2}(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\right\rangle-\frac{1}{4}\,\mathrm{Scal}\,|\varepsilon\cdot\zeta|^{2}-\frac{1}{18}\,\langle d\varphi\cdot(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\rangle$ (5.13)
$\quad+\frac{1}{162}\,\|\varphi\|^{2}|\varepsilon\cdot\zeta|^{2}-\langle F_{A}(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\rangle.$
From Proposition 5.5, if $\alpha\in\Omega^{1}(M,\mathrm{Ad}(P))$ is an infinitesimal deformation of the $\mathrm{G}_{2}$–instanton $A$, then $\varepsilon$ satisfies $D^{1/3,A}(\varepsilon\cdot\zeta)=-\frac{7}{12}\,\tau_{0}\,\varepsilon\cdot\zeta,$ hence
$\left\langle(D^{\frac{1}{3},A})^{2}(\varepsilon\cdot\zeta),\,\varepsilon\cdot\zeta\right\rangle=\frac{49}{144}\,\tau_{0}^{2}\,|\varepsilon\cdot\zeta|^{2}.$ (5.14)
From (2.23), we compute
$d\varphi$ $=2\left(X^{2356}+X^{2457}+X^{3467}\right)-\frac{3P_{1}}{2y_{2}}\left(X^{1237}-X^{1246}+X^{1345}-X^{1567}\right)$
$+\frac{3P_{3}}{2y_{2}}\left(X^{1234}-X^{1267}+X^{1357}-X^{1456}\right).$
where $P_{1}=a(a^{2}+b^{2})^{-1/2}$ and $P_{3}=b(a^{2}+b^{2})^{-1/2}$. Notice that $P_{1}^{2}+P_{3}^{2}=1$ and $P_{1}+P_{2}\leq 1$. On the other hand, from the proof of Theorem 2.1, we have $\tau_{0}=\frac{6}{7}\left(2+\frac{1}{y_{2}}\right)$ and therefore,
$\ast\tau_{3}$ $=d\varphi-\tau_{0}\psi$
$=$ $\left(\frac{26}{7}+\frac{6}{7y_{2}}\right)\left(X^{2356}+X^{2456}+X^{3467}\right)-\left(\frac{6}{7}+\frac{33}{14}\frac{P_{1}}{y_{2}}\right)\left(X^{1237}-X^{1246}+X^{1345}-X^{1567}\right)$
$+\left(\frac{6}{7}+\frac{33}{14}\frac{P_{3}}{y_{2}}\right)\left(X^{1234}-X^{1267}+X^{1357}-X^{1456}\right).$
Hence $\lvert\tau_{3}\rvert^{2}=3\left(\frac{26}{7}+\frac{6}{7y_{2}}\right)^{2}+4\left(\frac{6}{7}+\frac{33}{14}\frac{P_{1}}{y_{2}}\right)^{2}+4\left(\frac{6}{7}+\frac{33}{14}\frac{P_{3}}{y_{2}}\right)^{2}$ and using $P_{1}+P_{3}\leq P_{1}^{2}+P_{3}^{2}=1,$ we get
$\lvert\tau_{3}\rvert^{2}\leq\frac{4}{49}\left(241+276\,\frac{1}{y_{2}}+\frac{1125}{4}\,\frac{1}{y_{2}^{2}}\right).$ (5.15)
From (5.13), using (5.14)
$(*)$ $\leq\left(\frac{49\lvert\tau_{0}\rvert^{2}}{144}-\frac{\mathrm{Scal}}{4}-\frac{\lvert\tau_{0}\rvert}{18}+\frac{7}{162}+\frac{1}{18}\lvert\tau_{3}\rvert+\lvert F_{A}\rvert\right)\lvert\varepsilon\cdot\zeta\rvert^{2}.$
Thus, from (4.12) and (5.15), we obtain
$0$ $\leq\left(\frac{49\lvert\tau_{0}\rvert^{2}}{144}-\frac{\mathrm{Scal}}{4}-\frac{\lvert\tau_{0}\rvert}{8}+\frac{7}{162}+\frac{\lvert\tau_{3}\rvert}{18}+\lvert F_{A}\rvert\right)\lvert\varepsilon\rvert^{2}$ (5.16)
$\leq\Bigg[\frac{1}{4}\left(2+\frac{1}{y_{2}}\right)^{2}+3-\frac{6}{y_{2}}-\frac{3}{28}\left(2+\frac{1}{y_{2}}\right)+\frac{7}{162}+\frac{1}{63}\sqrt{241+\frac{276}{y_{2}}+\frac{1125}{4y_{2}^{2}}}+\frac{\sqrt{3}}{y_{2}}\Bigg]\lvert\varepsilon\rvert^{2}.$
Setting $x=\frac{1}{y_{2}},$ the right–hand side of (5.16) can be written as $f(x)\,|\varepsilon|^{2},$ where for $x\geq 0$
$f(x)$= $\frac{1}{4}(2+x)^{2}+3-6x-\frac{3}{28}(2+x)+\frac{7}{162}+\frac{1}{63}\sqrt{241+276x+\frac{1125}{4}x^{2}}+\sqrt{3}\,x$
The function $f$ is continuous on $0,\infty)$, moreover, a direct evaluation shows that $f(x)$ becomes negative on a bounded interval. In particular, $f(x)<0\quad\text{for all }x\in(\frac{5}{3},10).$ [^5] This interval is not maximal Equivalently, $f(y_{2})<0$ for all $y_{2}\in\left(\tfrac{1}{10},\tfrac{3}{5}\right)$. For such values of the metric parameter $y_{2}$, the right–hand side of the Weitzenböck inequality becomes strictly negative (see Figure 1), while the left–hand side is non–negative. This yields a contradiction unless $\varepsilon\equiv 0.$ Therefore, the instanton is rigid for this range of metric parameters.
xxf(x)f(x)
Figure 1: Plot of the right–hand side of the Weitzenböck inequality as a function of the metric parameter $y_{2}$. This shows that the expression becomes strictly negative for a non-empty interval of positive values of $y_{2}$.